OpenStudy (anonymous):

Simplify cosθ + cosθtan2θ. 1 cscθ secθ sin2θ

Nnesha (nnesha):

is it \[\tan^2 \theta ~?\]

Nnesha (nnesha):

\[ \huge\rm \cosθ + \cosθ\tan^2θ\] like this ?

OpenStudy (anonymous):

yes thats right

Nnesha (nnesha):

alright rewrite tan^2 in terms of cos and sin tan= ??

OpenStudy (anonymous):

what do u mean im a little confused

Nnesha (nnesha):

http://www.analyzemath.com/trigonometry/trigonometric_formulas.html bookmark this link you really need this for trig

Nnesha (nnesha):

i was asking what's the reciprocal of tan ?

OpenStudy (anonymous):

adj / opp

Nnesha (nnesha):

yes that's correct 

Nnesha (nnesha):

that's the relation between trig functions sin, cos , tan and csc ,sec ,cot are reciprocal of each other tan x=sinx over cosx

OpenStudy (anonymous):

ok

Nnesha (nnesha):

\[ \huge\rm \cosθ + \cosθ\color{ReD}{tan^2θ}\] we need to rewrite tan^2 in terims of sin cos \[\tan^2 \theta = \frac{\sin^2 \theta}{\cos^2 \theta}\]

OpenStudy (anonymous):

ok i got that now so wat do ido next

Nnesha (nnesha):

replace tan^2 with sin^2/cos^2 and then simplify \[ \huge\rm \cosθ + \cosθ*\color{ReD}{\frac{sin^2 \theta }{cos^2 \theta }}\]

Nnesha (nnesha):

cos^2 is same as cos times cos

OpenStudy (anonymous):

but how would i simplify it

Nnesha (nnesha):

\[ \huge\rm \cosθ + \cosθ*\color{ReD}{\frac{sin^2 \theta }{cos^2 \theta }}\] \[\cos \theta + \frac{\cos \theta *\sin^2 \theta}{\cos \theta *\cos \theta}\] cos^2 can be written as cos times cos now canyou cancel out anything ?

OpenStudy (anonymous):

oh ok so its the sin

Nnesha (nnesha):

sin ??

Nnesha (nnesha):

what do you mean ?

Nnesha (nnesha):